A 5.70 V battery is connected in series with a resistor and an inductor. The series circuit has a time constant of 589 μs, and the maximum current is 279 mA. What is the value of the inductance?
Answer #1

In Rl circuit the current in the resistor is maximum at initial .as the the time passes the current in the resistor decreases

I max = 589 us

Imax = V/R


R = V/Imax = 5.7/(279*10^-3) = 20.43 ohms

time onstant in RL circuit is given by t = L/R

L = R*t = 0.01203 Henry = 12.03 mH